what is the y intercept of the given linear equation?
y=3/2 x-1/4
A: (0, 1/4)
B: (0, -1/4)
C: (0, -4)
D: (0, 1)

Answer by Cromlix(4381)

Hi there,
With the equation y = 3/2x - 1/4
To find where the line cuts the 'y' axis
make 'x' = 0
y = 3/2x - 1/4
y = 3/2(0) - 1/4
y = -1/4
(0, -1/4) B
Hope this helps :-)